ITF found in Science Mode. From 6:07 UTC the Horizon started to drops severely ( see attached plot ), the DMS reported some anti-Squeezing Injected, so I contacted the related Oncall. We switched to LOW_NOISE_3 and the drops disappeared soon after. Back in Science Mode at 7:34 UTC.
The horizon drops reappeared again at 12:10 UTC, but this time no flags from the Squeezing resulted red. I contacted the ISC on call (D. Bersanetti ) to verify that the loops were working correctly, but no evident issues were found. I contacted again the SQZ Oncall and we switched back to LOW_NOISE_3, this seems a more stable condition. In agreement with the SQZ on-call, I left the ITF in Science Mode without Squeezing.
SUSP
Due to the strong wind of this morning, I switched to the related GIPC configuration from 10:03 UTC to 12:26 UTC.
Troubleshooting
From 6:53 UTC to 7:34 UTC and from 12:18 UTC to 13:03 UTC due to drops in the Horizon. The issue is still under investigation.
Guard Tour
5:27 UTC - 6:00 UTC;
7:25 UTC - 7 58 UTC;
9:31 UTC - 9:59 UTC;
11:01 UTC - 11:27 UTC.